OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4). y = four thirdsx + 8 y = four thirdsx + 3 y = −three fourthsx + 8 y = −three fourthsx + 3

OpenStudy (aravindg):

two lines are parallel means what they have in common ?

OpenStudy (anonymous):

since the requred eqn is parellel to the given line,slope of both the lines r equal

OpenStudy (aravindg):

it would have been better if she tried to answer that

OpenStudy (anonymous):

the slope of the above eqn is m=4/3(found out bybringing the above eqn to y=mx+c form)where m=slope....

OpenStudy (anonymous):

since the reqd eqn passes through -3,4 =x1,y1.....sub.in y-y1=m(x-x1)here m=4/3

OpenStudy (anonymous):

y-4=4/3(x+3)=>y=4x/3+8

OpenStudy (aravindg):

and thats what we call giving away answer ^^

OpenStudy (aravindg):

@ashishthomas7 violation of coc can result in suspension

OpenStudy (aravindg):

http://openstudy.com/code-of-conduct